Consider the following project: A new product requires an initial investment of $4000 and will be...

  1. Consider the following project: A new product requires an initial investment of $4000 and will be depreciated to an expected salvage of zero over 5 years. The price of the new product is expected to be $30, and the variable cost per unit is $10.The fixed cost is $1000the required return is 12%. What is the Accounting Break-even, what is the cash Break-even and what is the Financial Break even? Please show step by step, thank you!
Answer #1

Cash break even quantity = fixed cost / (price – variable cost)

Cash break even quantity =1000/ (30-10)= 1000/20

Cash break even quantity = 50

Accounting break even=(fixed cost + depreciation ) / (price – variable cost)

Depreciation = initial investment / life in years= 4000/5=800

Accounting break even=(1000+800) /(30-10) = 1800/20

Accounting break even= 90

Financial break even= fixed cost + operating cash flow /(price – variable cost)

Operating cash flow = initial investment / PV annuity factor

PV annuity factor= 1-1(1+r) ^-n / r

PVAF= 1-(1+0.12)^-5/.12 = 0.432573/0.12

PVAF= 3.6048

OCF = 4000/3.6048 = 1110

Financial break even=(1000+1110) / (30-10) = 2110/20

Financial break even=105.5